OpenStudy (anonymous):

For Medal : Help please with checking if i got the correct answer with simplifying complex fractions, question below. The answer i got is -1. Thanks :)

OpenStudy (anonymous):

\[\frac{ 35x ^{2} + 2x - 1 }{ 15x + 3 } \div \frac{ 2 - 98x ^{2} }{6x + 42x }\]

OpenStudy (anonymous):

Is that your answer? Or is that the question?

OpenStudy (anonymous):

that is the question, the answer i got was -1

jimthompson5910 (jim_thompson5910):

is that second denominator really 6x+42x?

OpenStudy (anonymous):

sorry, it is 6 + 42x, thank you

jimthompson5910 (jim_thompson5910):

the answer is not -1, try it again

OpenStudy (anonymous):

ok, thanks i will

jimthompson5910 (jim_thompson5910):

oh wait, i put in 6-42x my bad

jimthompson5910 (jim_thompson5910):

I changed it to 6+42x and it's saying -1 now, so you got it

OpenStudy (anonymous):

ok great, thanks you. do u have time to help with one more?

jimthompson5910 (jim_thompson5910):

sure, go for it

OpenStudy (anonymous):

one root of a quadratic equation is -3 + 2i. Determine the other root and write an equation that would result in these roots. I'm not sure how to do this one.

jimthompson5910 (jim_thompson5910):

if one root is a+bi, then the other root must be a-bi since all roots come in conjugate pairs

jimthompson5910 (jim_thompson5910):

so what's the conjugate of -3+2i?

OpenStudy (anonymous):

-3 - 2i

OpenStudy (anonymous):

how do u write the equation?

jimthompson5910 (jim_thompson5910):

well if the two roots are -3+2i and -3-2i, then we know that x = -3+2i or x = -3-2i x+3 = 2i or x+3 = -2i (x+3)^2 = (2i)^2 or (x+3)^2 = (-2i)^2 I'll let you finish

OpenStudy (anonymous):

why are the terms in the last line squared?

jimthompson5910 (jim_thompson5910):

well remember that \(\Large i = \sqrt{-1}\) if you square both sides, you get \(\Large i^2 = -1\)

jimthompson5910 (jim_thompson5910):

that's why I'm squaring both sides, to get rid of the imaginary i term

OpenStudy (anonymous):

thanks for explaining...so the answer would be x^2 + 6x + 11 or x^2 + 6x + 7?

jimthompson5910 (jim_thompson5910):

neither choice, here's why: x = -3+2i or x = -3-2i x+3 = 2i or x+3 = -2i (x+3)^2 = (2i)^2 or (x+3)^2 = (-2i)^2 (x+3)^2 = (2)^2*(i)^2 or (x+3)^2 = (-2)^2*(i)^2 (x+3)^2 = 4i^2 or (x+3)^2 = 4i^2 (x+3)^2 = 4i^2 (x+3)^2 = 4(-1) (x+3)^2 = -4 x^2 + 6x + 9 = -4 x^2 + 6x + 9 + 4 = 0 x^2 + 6x + 13 = 0

OpenStudy (anonymous):

sorry, i'm still confused...in line 4 how did u get 2^2 times i^2 from 2i^2?

OpenStudy (anonymous):

wouldn't it be 2 times i^2 and then -2 times i^2

jimthompson5910 (jim_thompson5910):

I used the rule that (ab)^c = a^c * b^c to go from (2i)^2 to (2)^2 * (i)^2

jimthompson5910 (jim_thompson5910):

it's NOT 2i^2 it's really (2i)^2

jimthompson5910 (jim_thompson5910):

the parenthesis make a big difference

OpenStudy (anonymous):

ok got it :) thanks so much for your help, i appreciate it!